OpenStudy (anonymous):

3b+9=-18 Easiest way to solve this?

OpenStudy (texaschic101):

1. subtract 9 from both sides 2. divide both sides by 3

OpenStudy (texaschic101):

its that simple :)

OpenStudy (anonymous):

Thank you! Also when i solve it I get a positive 9 even though I know its a negative 9. I don't know where i messed up.

OpenStudy (texaschic101):

3b + 9 = -18 -- subtract 9 from both sides 3b = -18 - 9 3b = - 27 -- divide both sides by 3 b = - 9